### v1

# Comment on Galina Weinstein’s article entitled “Convergences and Divergences: Einstein Poincaré and Special Relativity”

Abstract
In 2024, after thirty years of research on this subject, I published a book entitled: Poincaré, Einstein and the discovery of special relativity. An end to the controversy Ginoux2024 . In September 2025, Galina Weinstein published a review of this book entitled: Convergences and Divergences: Einstein Poincaré and Special Relativity (arXiv:2509.09361) in which she aimed “to show how Einstein’s ether-free, principle-based kinematics marked out a path that, unlike its contemporaries, became the canonical form of special relativity.” To this end, she claimed to reconstruct “in a novel way, the 1905 derivations of Einstein and Poincaré, highlighting their contrasting paths”. In her paper, she proposed a mathematical presentation of “the 1905 derivations of Poincaré and Einstein”. And then, she pretended to trace “their similarities and, more importantly, their differences.” Unfortunately, Weinstein’s review of my book contains a number of mistakes, falsehoods and misleading criticisms that I would like to point out here. To this aim, I will follow the structure of Weinstein’s paper and show section by section all the erroneous historical facts she has reported.
1 Introduction of Weinstein’s review
In her introduction, Galina Weinstein claims to reconstruct “in a novel way, the 1905 derivations of Einstein and Poincaré.” Unfortunately, all the mathematical derivations she pretends to “reconstruct” at subsection 2.4. and in sections 3. have been already presented in my book Ginoux2024 in more detailed way at Chapter 6 (her subsection 2.4) and in Miller’s contributions Miller1973 ; Miller1980 ; Miller1981 ; Miller2001 . Thus, her “novel way” dates back at least of 1973 and may be before.
2 Between Convention and Innovation of Weinstein’s review
2.1 The Elephant in the Room
In this section Galina Weinstein recalls:
“Ginoux also comments on Einstein’s 1955 letter to Carl Seelig, in which Einstein professed ignorance of Poincarés 1905 note [Poi05-1] and earlier Lorentz papers. He characterizes this as “surprising” in light of Einstein’s documented familiarity with contemporary literature, both in his published citations of the March 1905 quanta of light paper and in his role as a reviewer for the Beiblätter zu den Annalen der Physik [Gin]. While such claims bear on questions of influence rather than of strict priority, they invite broader historiographical reflection on the distinction between acquiring a mathematical toolkit and constructing a new conceptual architecture.”
I still claim that it’s very surprising that Einstein ignored the work of his predecessors in this article entitled “On Electrodynamics of moving bodies”, while in his seven other articles published between 1901 and 1905 in Annalen der Physik he cited all the most important work of his predecessors. Moreover, as a reviewer for the Beiblätter zu den Annalen der Physik, he was therefore well-versed in the scientific publication practices of his day.
To defend her point of view, Galina Weinstein cites Einstein’s 1955 letter to Carl Seelig in which Einstein explains that: “he knew Lorentz’s 1895 work but neither Lorentz’s later writings nor Poincaré’s related investigations, and that his 1905 work was “in this sense” independent.”
Here, the conflict of interest is obvious since the only witness on which these claims are based is Einstein himself. So, we must believe him according to Galina Weinstein because Einstein necessarily always tells the truth. This is unfortunately not the case and the biographies (see Albert Einstein Demystified, Ginoux Ginoux2020 ) I wrote on Einstein demonstrate this. Indeed, Einstein, like many others, lied to his wife, his children and also to his colleagues. So, why should we believe what he says about this article? Sorry but this is clearly not enough.
Then, Galina Weinstein wrote:
“What was new, he emphasized, was recognizing that the Lorentz transformation applied beyond electrodynamics, reaching the general structure of space and time, and that Lorentz invariance was a universal constraint on physical theories.”
In fact, the original Lorentz transformation of 1904 was uncomplete and so, was not invariant. In May 1905, one month before Einstein has submitted his article to Annalen der Physik, Poincaré had already stated a complete transformation to which he gave the name of Lorentz and for which he proved its invariance. Moreover, it was not until 1910 that Einstein gave to the transformation (that he has established more than one month after Poincaré) the name of Lorentz. Thus, Einstein established on June 30, 1905 a transformation which is perfectly identical to that Poincaré stated at least one month before and which is not the original Lorentz transformation.
In the following of her subsection 2.1, Galina Weinstein makes reference to Gerald Holton and explains that:
“The omission of citations to either Poincaré or Lorentz’s 1904 paper, he suggests, is best read in light of Einstein’s normal practice of acknowledging sources he actively used; in the very same paper, Einstein twice names Lorentz when referring to the electron theory as presented in the 1895 Versuch, which he had read [Ein05, Hol60].”
How could Holton know wether or not Einstein had actually used the articles that he quoted? Many searchers quote some papers they have not even read and many searchers don’t quote for many reasons some papers they have read. This argument is simply unacceptable.
Then, Galina Weinstein writes that:
“Where Lorentz (and Poincaré) began from the transformations as a given, Einstein deduced them from two postulates - the relativity principle and the constancy of the speed of light - thus arriving by a distinct route [Hol60].”
This sentence leads to a question concerning the development of the theory of special relativity.
How and why Lorentz was led to develop his transformation of 1904?
He used the principle of relativity formalized by Poincaré at the Saint Louis congress in 1904 and according to which “the laws of physical phenomena must be the same for a stationary observer as for an observer carried along in a uniform motion of translation” Poin1904 to show that Maxwell’s equations of electromagnetism are invariant. Of course, Lorentz and Poincaré began from the transformation (without s) to prove its invariance according to the relativity principle. This is confirmed by Lorentz himself who wrote in 1921 (nearly ten years after Poincaré’s death):
“I did not succeed in obtaining the exact invariance of the equations (…).
I did not establish the principle of relativity as rigorously and universally true. Poincaré, on the contrary, obtained a perfect invariance of the equations of electrodynamics, and he formulated the ’postulate of relativity’, terms which he was the first to employ (…) Lorentz1921 .”
Then, Galina Weinstein explains:
“In Poincaré’s presentation, simultaneity, time measurement, and the operational meaning of coordinates remained within the conceptual boundaries of the ether theory.”
Here, Galina Weinstein refers to Poincaré’s article entitled “La mesure du temps” published in 1898 Poin1898 . Unfortunately, this contribution of Poincaré does not contain the word ether! Galian Weinstein discusses the question of ether in subsection 2.5 we analyze below.
2.2 Between Bern and Paris: No Telegraph
In this subsection Galina Weinstein writes:
“At that time, the 25-year-old patent examiner in Bern stood outside the scholarly correspondence networks through which such material typically circulated and had not yet met Lorentz. Surviving documentation records no communication from Lorentz to Einstein in this period, and Einstein’s first known exchange with a leading academic - his correspondence with Max Planck - dates from roughly a year later.”
This fairy tale reported from year to year by some historians of science is a pure fiction. As I have explained in my book Ginoux2024 , the links are actually quite numerous. First of all, Planck was associate editor of the journal Annalen der Physik, in which Einstein published his first articles as early as 1901. Indeed, his famous article “On the electrodynamics of moving bodies,” considered the founding text of the special theory of relativity, was in fact his eighth publication in this journal. In addition, from 1905 onwards, Einstein wrote reviews of articles published in other international journals for the Beïblätter zu den Annalen der Physik, i.e., the supplement to Annalen der Physik (see Chap. 7 of my book). Moreover, according to Klein and Needell Klein , Einstein reviewed a work by Planck published in 1906. An analysis of Einstein’s correspondence shows that their first epistolary exchanges date back to the 6th July 1907. This is absolutely impossible, for several obvious reasons. Firstly, since Einstein worked for the supplement of Annalen der Physik, he must have had contact with its editors, at least for signing his employment contract or for sending back his article reviews. Moreover, Einstein was still looking for an academic position at the university, as evidenced by this letter to his girlfriend Mileva Maric dated the 4th April 1901:
“Soon I will have honored all physicists from the North Sea to the southern tip of Italy with my offer!EinCorr1 .”
It is therefore astonishing that Einstein did not send his application to Planck. Finally, in a letter from Einstein to his friend Maurice Solovine dated the 27th April 1906, we read:
“My papers are much appreciated and are giving rise to further investigations. Professor Planck (Berlin) has recently written to me about that EinCorr5 .”
It is clear from this letter that Planck and Einstein had already corresponded prior to the 27th April 1906. Unfortunately, these letters have disappeared.
At the end of this subsection, Galina Weinstein writes the following sentence that is the leitmotif of her article:
“Poincaré’s synchronization and Einstein’s synchronization may look similar at the procedural level, but they are embedded in fundamentally different conceptual frameworks.”
and that we could summarize as follows: Same but Different.
2.3 Priority Thread, in One Breath
In this subsection Galina Weinstein writes:
“Ginoux’s book adopts a formalist, sequence-oriented historiography, in which the systematic collation of equations, dates, and correspondence is used to reconstruct the relative timing and scope of contributions. On this basis, he attributes to Poincaré, by May-June 1905, a body of results encompassing the corrections to Lorentz’s 1904 formulas, the symmetric transformation form with $l=1$, the group property, and the relativistic velocity-addition law as presented in the June 5 Comptes rendus note [Gin]. For Ginoux, these achievements, combined with Poincaré’s articulation of the relativity principle, constitute the formal underpinnings of special relativity.”
Let us recall here that this not Ginoux who considered that these achievements “constitute the formal underpinnings of special relativity” but Einstein himself. In fact, Einstein wrote in 1935 that he considered:
“the Lorentz transformation [as] the real basis of the special relativity theory Einstein1935 .”
At the end of this subsection, Galina Weinstein repeats again:
“Verbal correspondences, such as the parallel between the title of Einstein’s paper and a phrase from Poincaré’s 1904 Saint Louis lecture [Poi04], are noted alongside recurrent juxtapositions of formulations that are often described as “identical” in substance, despite differences in expression.”
As recalled in my book Ginoux2024 , in 1904 at Saint Louis, Poincaré concluded his oral presentation by this sentence:
“It is a question before all of endeavoring to obtain a more satisfactory theory of the electrodynamics of moving bodies Poin1904 .”
This expression is very interesting because it is exactly the title of Albert Einstein’s article, “Zur Elektrodynamik bewegter Körper” (On the electrodynamics of moving bodies), Annalen der Physik, 17(10), 891-921, received the 30th June 1905 and published the 26th September 1905.
Thus, is there any difference? Absolutely none. This doesn’t prove that Einstein has read Poincaré’s contribution of 1904. This is just a striking coincidence and there are many more.
2.4 The Ghost Prefactor
This is probably the most surprising part of Weinstein’s article. Concerning the following factor introduced by Einstein in his article:
$\varphi(v)=a(v)\dfrac{1}{\sqrt{1-\frac{v^{2}}{c^{2}}}}$ (1)
she wrote:
“Ginoux treats (1) as a purposeful nudge toward the Lorentz transformation rather than a neutral reparametrization..”
First of all, it is not only Ginoux who “treats (1) as …” but also Professor Arthur I. Miller. Then, it is unclear how and why Einstein introduced this factor as recalled by Miller who wrote in 1981 Miller1981 :
“How did Einstein know that he had to make the further substitutiona equals $a=\varphi(v)\sqrt{1-\frac{v^{2}}{c^{2}}}$ in order to arrive at those space and time transformations in agreement with the postulates of relativity theory?”
and he added:
“But why did Einstein make this replacement. It seems as if he knew beforehand the correct form of the set of relativistic transformations Miller1981 .”
Unfortunately, Galina Weinstein seems to be unable to answer to these questions. At the end of this subsection Galina Weinstein makes a new mistake by writing:
“Using the final Lorentz transformation (38), after fixing $a(v)$ and $\varphi(v)=1$, he obtained the relativistic addition law [Ein05].”
Contrary to what has written Galina Weinstein, in his original paper, Einstein first used a long and tedious computation to obtain the relativistic addition law (see my the chapter 7 of my book Ginoux2024 ). In Fig. 1 below, I have reproduced a screen shot of Einstein’s original paper Einstein1905 concerning his proof of the relativistic addition law
Figure 1: Screen shot of Einstein’s original paper
As one can see in Fig. 1, contrary to what claims Galina Weinstein, this is not by using the final Lorentz transformation that Einstein first obtained the relativistic addition law.
At the end of this subsection, Galina Weinstein explains:
“Ginoux suggests that the paucity of citations in 1905 indicates dependence on Lorentz / Poincaré. Regardless of historical editorial practice, that claim is orthogonal to the logic of the derivation.”
Here again, Galina Weinstein is in the interpretation of what I have written. I have just explained in my book that it’s important to demonstrate that the argument too often put forward by some historians of science that “it wasn’t customary to cite one’s sources in 1905”(as it seems to be the case for Galina Weinstein here) doesn’t hold water. In fact, three months before the publication of his article “On the electrodynamics of moving bodies”, Einstein published an article entitled “On a heuristic point of view concerning the production and transformation of light” Einstein1905a on March 18, 1905, also in Annalen der Physik. In this work on the explanation of the photoelectric effect, for which he was awarded the Nobel Prize in Physics in 1922 (for the year 1921), Einstein made no fewer than seven quotations.
So why does he quote his peers in March and not in June? Once again, this fact raises questions.
2.5 Poincaré’s Ether vs. Einstein’s Ether
In this subsection Galina Weinstein discussed the existence or non-existence of ether which would be a decisive step in the development of the theory of special relativity theory. Here again, this is not the good question to be addressed. Of course Poincaré has kept the luminiferous ether as many other scientists at that time, i.e. like Lorentz and Planck and before them James Clerk Maxwell. At that time, it was commonly admitted that light was propagating in a medium called luminiferous ether. Contrary to what claims Galina Weinstein, Einstein did not abolish ether. In his article, Einstein Einstein1905 exactly said: “Die Einführung eines “Lichtäthers” wird sich insofern als überflüssig erweisen” (The introduction of a “light ether” will prove superfluous in this respect). The word “überflüssig (superfluous)” does not mean that he abolished ether but that he didn’t need to use it as it was nearly the case. In fact as highlighted in my book, Lorentz and Poincaré were working on a problem of contraction of electron moving at velocity close to that of light. Thus, they tried to explain such a contraction by means of force and more particularly the Lorentz electromagnetic force. Thus, their approach was dynamic (from the Greek dyn which means force). The title of Poincaré’s two main contributions entitled “Sur la dynamique de l’électron” (On the dynamics of the electron) Poin1905 ; Poin1906 is enough to prove that. The reason why Poincaré kept the luminiferous ether is based on his dynamic approach as highlighted by this sentence he wrote in 1900:
“If we did not wish to change the whole of the science of mechanics, we should have to introduce the ether, in order that the action which matter apparently undergoes should be counterbalanced by the reaction of matter on something. Poin1900a .”
Poincaré justified the existence of a luminiferous ether as a convenient hypothesis, explaining that without it, Newton’s third law - the action-reaction principle - would no longer be respected. Thus, his dynamic approach consisted in extending Newton’s classical mechanics to Maxwell’s electromagnetism in what he called a New Mechanics (Mécanique Nouvelle). By completing the latest transformation provided by Lorentz in 1904, Poincaré proved that the resulting transformation he had thus obtained formed a group of invariance of the Dynamics.
Einstein’s article Einstein1905 shares into two main sections entitled kinematic and dynamic part. As recalled by the title of the first part, his approach is, first of all, kinematic. Thus, his aim is completely different from that of Lorentz and Poincaré. At first, he is not interested in the contraction of an electron moving at a velocity close to that of light. The problem to which he is faced is the Doppler-Fizeau effect. This phenomenon can easily be observed in our everyday lives when an emergency vehicle passes by. When the vehicle approaches, the sound produced by its siren seems higher-pitched, whereas it seems lower-pitched when it moves away. The question for Einstein was then to know what would happen if we replace the acoustic or mechanical wave, i.e. the siren, with an electromagnetic wave, i.e. light. To this aim, Einstein used this famous metaphor: if a person travels at a speed close to that of light, will he be able to see his face in a mirror placed in the direction of his travel? Following this idea, Einstein analyzes motion independently of the causes (forces) that produce it. More precisely, Einstein studied the propagation of a light signal from one frame of reference to another. However, as recalled above, normally Einstein should have kept the luminiferous ether since at that time, it was commonly admitted that light propagates in such a medium. But he didn’t say anything about the medium in which his light signal propagated. It’s only in the second part of his article that he considers an empty space, without giving any definition. By using his kinematic approach, he was able to establish a transformation which is exactly that given by Poincaré in May 1905 and published by Poincaré Poin1905 on June 5th 1905 in the Comptes Rendus, thus at least three weeks before Einstein had submitted his article.
At the end of this subsection, Galina Weinstein explains:
“When Einstein did speak of an ether, it was in a sense that differed fundamentally from Lorentz’s construct.”
To support it, Galina Weinstein repeats the sentence uttered at the time by John Stachel, whom she thanks at the end of her article, as if repeating an argument could give it a truth value:
“the ether he reintroduced differed fundamentally from the ether he had banished [Sta-01].”
But to defend the indefensible, Galina Weinstein will now provide us with the following weighty argument.
“In a 1916 letter to Lorentz, Einstein himself, half-diplomatically and half-seriously, had remarked that general relativity was “closer to the ether hypothesis” than special relativity [CPAE8], Doc. 222. The remark was a gesture of respect: Lorentz still clung to ether, and Einstein, who revered him, framed his own theory in language that Lorentz could recognize. The Leiden lecture was thus homage as much as physics - a theatrical bow to Lorentz, even as the stage and script had already shifted.”
Thus, according to Galina Weinstein, when Einstein reintroduced the ether in a lecture given at Leiden, this is just “a theatrical bow to Lorentz” and we should not consider that he really reintroduced the luminiferous ether. Let us recall below the conclusion of Einstein’s lecture in Leiden Einstein1921 which has been subtly omitted by Galina Weinstein:
“Recapitulating, we may say that according to the general theory of relativity space is endowed with physical qualities; in this sense, therefore, there exists an ether. According to the general theory of relativity space without ether is unthinkable; for in such space there not only would be no propagation of light, but also no possibility of existence for standards of space and time (measuring-rods and clocks), nor therefore any space-time intervals in the physical sense. But this ether may not be thought of as endowed with the quality characteristic of ponderable media, as consisting of parts which may be tracked through time. The idea of motion may not be applied to it Einstein1921 .”
This quotation is very interesting because it proves that in his 1905 paper, Einstein Einstein1905 should not have considered the luminiferous ether as superfluous because, as he wrote in his conclusion above, in a space without ether, propagation of light would be impossible. Moreover, when Einstein has decided to generalize the special relativity theory, i.e. to analyze the gravitational force able to bent light, he was obliged to reintroduce an ether as Poincaré did before him to analyze in a dynamic approach the contraction of an electron moving at velocity close to that of light. Thus, the argument of Galina Weinstein according to which the Leiden lecture was “a theatrical bow to Lorentz” is simply irrelevant. To confirm that Einstein had really introduced ether in the framework of general relativity theory, let us recall that in January 1920 Einstein had written a remarkable article for the journal Nature, which was never published and which he probably used for his lecture in Leiden. In his conclusion, Einstein wrote:
“Again, empty space seems to be endowed with physical properties, that is, not physically empty as it appeared in the special theory of relativity. Therefore, one can say the ether has been resurrected in the theory of general relativity, even though in a (newer) more sublime form. The ether of the general theory of relativity differs from the one in old optics by not being a substance in the sense of mechanics. Not even the concept of motion can be applied to it EinCorr7 .”
2.6 The Light Postulate
In this subsection Galina Weinstein explains:
“Ginoux regards Einstein’s 1905 light postulate as curious, since it appears to place the invariance of $c$ at the origin of the relativity principle, rather than the other way around.”
I still claim that Einstein’s 1905 light postulate is curious, since in my opinion it should have been like for Poincaré a consequence of the relativity principle and not the contrary. From the relativity principle, one can deduce the classical addition-law which is no more valid for light as proven by Michelson-Morley experiments. So, Einstein is perfectly right when he explains that:
“We will raise this conjecture (the purport of which will hereafter be called the Principle of Relativity) to the status of a postulate, and also introduce another postulate, which is only apparently irreconcilable with the former, namely, that light is always propagated in empty space with a definite velocity $c$ which is independent of the state of motion of the emitting body. Einstein1905 .”
However, in my opinion, the invariance of velocity of light is not a postulate irreconcilable with the principle of relativity but a consequence of this principle. That’s the reason why I invoked the works of Pr. Jean-Marc Lévy-Leblond LevyLeblond1976 .
In this subsection Galina Weinstein writes:
“The identification of $c$ as the limiting speed remains an essential, physically motivated step - one that Einstein’s 1905 formulation incorporates from the outset with full conceptual economy.”
Let us recall that at Saint Louis, in September 1904, Poincaré had already stated the principle of invariance of the velocity of light:
“From all these results, if they were confirmed, would arise an entirely new mechanics, which would be, above all, characterized by this fact, that no velocity could surpass that of light Poin1904 .”
In May 1905, in one of the letters written to Lorentz (see Fig. 2), Poincaré explained that in his demonstration, aimed at completing the Lorentz transformation, he has chosen the units such that of light $c=1$. Obviously, if Poincaré posed $c=1$, it is because he considered the velocity of light was the same in all reference frames.
3 Einstein’s and Poincaré’s Derivations of Weinstein’s review
3.1 Poincaré’s May 1905 Letters to Lorentz
In this subsection, Galina Weinstein reproduces parts of what I have published in Chapters 5 & 6 of my book Ginoux2024 . Then, she recalls that by applying successively two transformations, Poincaré stated the following formula:
$\varepsilon^{\prime\prime}=\dfrac{\varepsilon+\varepsilon^{\prime}}{1+\varepsilon\varepsilon^{\prime}}$ (2)
Then, Galina Weinstein explains:
“Why is equation (26) (Eq. (2) above) not the velocity addition law? The symbols $\varepsilon$, $\varepsilon^{\prime}$, $\varepsilon^{\prime\prime}$ are group parameters labeling Lorentz transformations (essentially, dimensionless rapidities). Equation (26) is therefore the group composition law for successive boosts, not a physical law of how material particle velocities add.”
Unfortunately, this argument is again irrelevant because Poincaré wrote in his letters to Lorentz reproduced in Figs. 2 & 3 below.
“Let $\varepsilon$ be the speed of translation with that of light taken as unity.”
It follows that the symbols $\varepsilon$, $\varepsilon^{\prime}$, $\varepsilon^{\prime\prime}$ are not group parameters but “speed of translation” according to Poincaré himself. Thus, the formula established by Poincaré in his May 1905 letters to Lorentz (see Figs. 2 & 3) is indeed the relativistic velocity addition law.
Figure 2: Poincaré’s May 1905 original letter to Lorentz, by courtesy of Noord-Hollands Archief, [Fonds Poincaré], NHA-9423.
Figure 3: Poincaré’s May 1905 original letter to Lorentz, by courtesy of Noord-Hollands Archief, [Fonds Poincaré], NHA-9423
3.2 Poincaré’s 1906 Derivation in the Rendiconti paper
In this subsection, Galina Weinstein reproduces parts of what I have published in Chapters 5 & 6 of my book Ginoux2024 . From her computations she obtains the following equation:
$\xi^{\prime}=\dfrac{\xi+\varepsilon}{1+\varepsilon\xi}$ (3)
Then, she explains in the previous subsection that:
“The actual velocity transformations (36) (Eq. (3) above) require differentiating the Lorentz transformation (24) as Poincaré himself later derived in the Rendiconti di Palermo memoir.”
In the subsection 3.2 she confirms that her equation (36) (Eq. (3) above) is indeed the relativistic velocity addition law because in his Rendiconti di Palermo memoir Poincaré has used a differential operator to derive it. This is simply incredible! The first argument of Galina Weinstein presented in her subsection 3.1 was that Poincaré used group symbols. But here, $\xi$ and $\varepsilon$ are still dimensionless group symbols. So, according to her own argument, her equation (36) (Eq. (3) above) should not be considered as the relativistic velocity addition law.
Could Galina Weinstein explain us the difference between her equations (36) and (26) (Eq. (3) and Eq. (2) above).
There are indeed two different ways to derive the relativistic velocity addition law. The first consists in proving that the Lorentz transformation is invariant. This is what Poincaré has done in his May 1905 letters to Lorentz (see Figs. 2 & 3). The second consists in using the differential operator. This is what Poincaré has done in his Rendiconti di Palermo memoir but not Einstein in the first part of his article Einstein1905 (see Fig. 1). Let us notice that if the last method does indeed provide the relativistic velocity addition law, it does not prove the invariance of the Lorentz transformation and therefore does not allow to demonstrate that it forms an invariance group of the Dynamics. Concerning Einstein, let us notice that he didn’t prove that the Lorentz transformation forms an invariance group of the Dynamics. He wrote in his paper:
“…such parallel transformations - necessarily - form a group. Einstein1905
Moreover, this sentence without any proof is very surprising since at that time, Einstein had never spoken about the group theory neither in his publications nor in his correspondence while Poincaré had already been in contact several times with Lazarus Fusch and Felix Klein and had already published many papers on this subject. May be another striking coincidence for Einstein?
3.3 Einstein’s 1905 Derivation of The velocity and Charge Density Transformation
In this subsection, Galina Weinstein presents Einstein’s long and tedious computation of the velocity and charge density already presented in the chapter 7 of my book Ginoux2024 . This is also very important to recall that according to Einstein and so to Galina Weinstein, Einstein had never been aware of Lorentz 1904 paper Lorentz1904 in which he gave some erroneous expressions of the velocity and charge density. This has been confirmed by Lorentz in 1921. He wrote:
“Poincaré, on the contrary, obtained a perfect invariance of the equations of electrodynamics, and he formulated the postulate of relativity, terms which he was the first to employ. Indeed, stating from the point of view that I had missed, he found the formula ([of the velocity and charge density]). Let us add that by correcting the imperfections of my work he never reproached me for them Lorentz1921 .”
How could Einstein know that he had succeeded to obtain the correct expressions for the velocity and density of charge without reading Lorentz 1904 paper? Another coincidence probably? Let us notice that Einstein’s expressions of the velocity and density of charge are the same as those obtained in May-June by Poincaré, as highlighted in the chapter 7 of my book Ginoux2024 ,
3.4 The Addition Law that Made the Difference
In this subsection, Galina Weinstein explains that:
“Einstein did not, and in 1905 could not, adopt this purely formal route [group theory], since his method was heuristic in character.”
This surprising to read that Einstein has the right to use a heuristic method to obtain the relativistic velocity addition law while Poincaré cannot use such a method to keep the luminiferous ether.
4 In the Shadow of Light Beams of Weinstein’s review
4.1 The Relativity of Recognition
In this subsection, Galina Weinstein recalls that:
“Ginoux’s examination of the Nobel dossiers casts the episode less as a mystery of merit than as a study in how institutions quietly decide what will be remembered. Despite repeated nominations, Poincaré’s case never advanced. The committee preferred tangible experiment to abstract theory, mistrusted mathematical style in physics, and managed national sensibilities with a cautious hand. The reports are respectful, sometimes admiring, but never decisive. His work is acknowledged in passing, yet consistently absorbed into Lorentz’s theory or filed under collective progress. By the time of his death in 1912, the record had settled into a pattern of praise without credit - an archive of tributes that effectively erased their subject.”
Galina Weinstein has just omitted again two crucial points. The first is that starting from 1910, Einstein gave to the transformation he obtained in 1905 the name of Lorentz. Einstein wrote:
“These transformation equations were introduced in a very successful manner into electro-dynamics by M. Lorentz. We’ll refer to them as the Lorentz transformation Einstein1910 .”
Why didn’t Einstein do it before? Probably another coincidence. The most probable reason is that if he had done that in 1905 it will have proven that he had read Poincaré. The second crucial point is that in 1912 one discovers that Wien nominated Lorentz and Einstein for their common discovery of the special theory of relativity. Lorentz’s attitude is inexplicable to me. In May 1905, he writes to Poincaré to ask him his help concerning the transformation he had obtained gropingly Lorentz1921 , Poincaré send him back the complete transformation and the correct expressions for the velocity and density of charge. And a few years later instead of supporting him for the Nobel Prize of Physics (as Poincaré did for him in 1901) he applies for a second Nobel Prize with Einstein. This is very surprising even shocking. But there are probably some unknown reasons for this turnaround.
4.2 La problématique: Why Did Poincaré Not Claim Authorship of Special Relativity?
In this subsection, Galina Weinstein explains that:
“And yet, a few years later, in his 1912 London lecture, he was speaking the idiom of Minkowski space and edging toward Einstein’s own perspective (…) One even hears the unmistakable echo of Minkowski in his remark that the fourth coordinate is best taken as $\sqrt{-1}$.”
This is completely untruth. In his Rendiconti di Palermo memoir Poin1906 , Poincaré introduced two years before Minkowski Minkowski1908 the concept of four-vector as highlighted by the page 168 of his famous memoir (see Fig. 4 below).
Figure 4: Page 168 of Poincaré’s Rendiconti di Palermo memoir Poin1906 .
Then, Galina Weinstein explains that:
“The irony is hard to miss: Ginoux asks why Poincaré never claimed authorship of relativity [Gin], yet Poincaré’s own words, in the last public address of his life, sound less like a claimant and more like a convert. The revolution, it seems, was already underway - only the naming rights remained unspoken.”
This is again false. Poincaré gave his last lecture at the École Supérieur des Postes et Télégraphes today SupTelecom Paris in July 1912 a few days before his death as confirmed by the subtitle Poin1912 .
5 Conclusion: Saving Private Einstein
It is very surprising to observe for decades how some historians of science are able not to reconstruct but really to rewrite the development of the theory of special relativity by inventing imaginary facts or by interpreting real facts in a incredible manner as it is the case for the article of Galina Weinstein. First of all she explains that she will “reconstruct in a novel way the 1905 derivations of Einstein and Poincaré” although she had only recopied some results already published in my book Ginoux2024 and in the contributions of Pr. Miller Miller1973 ; Miller1980 ; Miller1981 ; Miller2001 . Thus, her “novel way” dates back at least of 1973 and may be before. Her aim is clearly to show that even if Poincaré has published the main results concerning the theory of special relativity (complete Lorentz transformation and its invariance while using group theory, relativistic velocity addition law, …), Einstein must be considered as the unique father of this theory.
Then, Galina Weinstein uses different types of arguments to support her point of view. The first one is the “truth”. As an example, Einstein claimed several times that he had never read Poincaré’s note and memoir, so it must be true and we must believe in his words even if he has lied in several circumstances to his wives, children and colleagues. If it is not enough, she invokes Gerald Holton according to whom Einstein only quoted papers he had read.
The second type of argument is “the selection of facts” which thus makes it possible to rewrite a history which corresponds to the thesis that one wishes to defend. As an example Galina Weinstein omitted to recall that the original Lorentz transformation of 1904 was uncomplete and that Lorentz contacted Poincaré in May 1905 to ask him to solve this problem.
The third type of argument is more dangerous but very classical: “the falsehoods”. As an example, she explains that in his contribution entitled “La mesure du temps” Poincaré Poin1898 already used the concept of luminiferous ether although even the expression does not appear in this paper as it is easy to verify. Then, she tries to make us believe that Einstein who was working for Annalen der Physik since 1901 and who has published before 1905 eight articles in this revue had never been in contact with one of the editor in chief Max Planck.
Unfortunately for Einstein’s defenders, the date of the publication of Poincaré’s contributions as well as that of Einstein is very well-known and well-documented: on June 5th 1905, Poincaré published his note Poin1905 and on June 30th 1905, Einstein submitted his article to Annalen der Physik Einstein1905 and so, it is obvious that Einstein 1905 famous paper is therefore at least several weeks later.
Here appears a new kind of argument: “same but different”. Poincaré and Einstein’s results seem to be the same but they are different. Why? This is not crystal clear. Thus, in her article Galina Weinstein made use of these four kinds of arguments to support her point of view which can be summarized as follows. First, she is obliged to admit that Einstein’s famous paper is subsequent to Poincaré’s contributions. So, she explains that contrary to Einstein, Poincaré kept the luminiferous ether and this fact precludes to consider that he could have laid the foundations of the theory of special relativity. However, Galina Weintein is also obliged to recognize that during his Leiden lecture in 1920 before Lorentz, Einstein claimed that, in the framework of the theory of general relativity, “there exists an ether”. Faced to such Einstein’s inconsistency (this is not the only one), she explained that when Einstein reintroduced the ether in a lecture given at Leiden, this is just “a theatrical bow to Lorentz” and we should not consider that he really reintroduced the luminiferous ether. Such an argument is simply irrelevant. But the most incredible is her argument to explain that Poincaré cannot have derived the relativistic velocity addition law while using the invariance of group because the symbols involved are those of group parameters. In her next subsection she recognizes that Poincaré has finally derived the relativistic velocity addition law while using a differential operator applied to a transformation (that Poincaré has completed and to which he has given the name of Lorentz) involving the same group parameters symbols. Curiously, Galina Weinstein made no comment on the fact that Einstein provided no proof that the Lorentz transformation forms a group of invariance of the Dynamics. She also omitted to recall that Lorentz expressions for the velocity and charge density were erroneous. A fact that Einstein was supposed to ignore since he was (according to him) unaware of Lorenz 1904 publication. In her article, Galina Weinstein also omitted to recall that Lorentz and Einstein applied in 1912 for the Nobel Prize of Physics for their common discovery of the theory of special relativity. A fact that raises questions. Her last subsection is also for Galina Weinstein the occasion to present a fake news by claiming that Poincaré has made use of Minkowski’s fourth-vector although Poincaré introduced it two years before him in his Rendiconti di Palermo memoir Poin1906 .
It is very regrettable that some historians of science are capable of using such methods to defend an indefensible point of view which does not stand up to analysis of the facts. Let us recall to them Poincaré’s words:
“Thinking must never submit itself, neither to a dogma, nor to a party, nor to a passion, nor to an interest, nor to a preconceived idea, nor to whatever it may be, if not to facts themselves, because, for it, to submit would be to cease to be Poin1909 . ”